OpenStudy (anonymous):

A train starts from rest and travels for 5.0 s with a uniform acceleration of 1.5 m/s2. What is the final velocity of the train?

OpenStudy (anonymous):

\[v_f=v_i+a \times t\] \[v_f = 0 + 1.5 \times 5 = 7.5_{m/s}\]

OpenStudy (anonymous):

Thanks